OpenStudy (anonymous):

A washer and a dryer cost $450 combined. The cost of the washer is two times the cost of the dryer. What is the cost of the dryer?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Alright so almost the same one here...washer (w) and dryer (d) combined would total 450 so w + d = 450 we also know that the washer is 2 times the cost of the dryer...so w = 2d Another nice substitution :)

OpenStudy (johnweldon1993):

Can you take it from there?

OpenStudy (anonymous):

Uhmm half of 450 is 225 soo

OpenStudy (johnweldon1993):

So we have w + d = 450 w = 2d 2d + d = 450 3d = 450 d = 450/3 d = ?

OpenStudy (anonymous):

soo the washer is 300 and the dryer is 150 because 300 is times of 150

OpenStudy (johnweldon1993):

That would be correct :)